OpenStudy (anonymous):

Use the elimination method to solve the system. 6x + 3y = –12 6x + 2y = –4 (10, –16) (2, –8) (–2, 8) (–10, 16)

OpenStudy (anonymous):

@blurbendy is this like the other one

OpenStudy (anonymous):

solve for y

OpenStudy (blurbendy):

not really. we used substitution the last time. they want you to use the elimination method, so basically you have to choose which variable you want to get rid of first. i would recommend multiplying the entire second equation by -1

OpenStudy (anonymous):

kk

OpenStudy (blurbendy):

what do you get?

OpenStudy (anonymous):

-6x-2y=4

OpenStudy (blurbendy):

okay, so now you can add the two equations together and eliminate the x 6x + 3y = -12 + -6x -2y = 4

OpenStudy (blurbendy):

what does your new equation become?

OpenStudy (anonymous):

y=-8

OpenStudy (blurbendy):

perfect, so now substitute -8 for y in either the top or bottom equation. your choice.

OpenStudy (anonymous):

kk

OpenStudy (blurbendy):

tell me what you get

OpenStudy (anonymous):

@Ashley_Cute This method is very simple you just need to cancel one variable that can be x or can be y 6x + 3y = –12 6x + 2y = –4 - - + ======== y=-8 Put the value od y in any equation 6x + 3y = –12 6x+3(-8)=-12 6x-24=-12 6x=-12+24 6x=12 x=12/6 x=2

OpenStudy (anonymous):

x=-2

OpenStudy (blurbendy):

which equation did you substitute -8 in for?

OpenStudy (anonymous):

+2 not -2

OpenStudy (anonymous):

lol thx but i want learn how to do it not the answer

OpenStudy (anonymous):

I explained it :)

OpenStudy (anonymous):

well it shows me so thx

OpenStudy (anonymous):

xD

OpenStudy (blurbendy):

which equation did you substitute -8 in for?

OpenStudy (anonymous):

the top one

OpenStudy (blurbendy):

so you'd end up with 6x + 3(-8) = -12 6x - 24 = -12 6x = 12 x = 2 You must've got one of your signs switched around somewhere

OpenStudy (anonymous):

so its 2 not -2

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

thx

OpenStudy (blurbendy):

yw

OpenStudy (anonymous):

Yw :)